### v2

# LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models

Abstract
The emergent reasoning and Theory of Mind (ToM) abilities demonstrated by Large Language Models (LLMs) make them promising candidates for developing coordination agents. In this study, we introduce a new LLM-Coordination Benchmark aimed at a detailed analysis of LLMs within the context of Pure Coordination Games, where participating agents need to cooperate for the most gain. This benchmark evaluates LLMs through two distinct tasks: (1) Agentic Coordination, where LLMs act as proactive participants for cooperation in 4 pure coordination games; (2) Coordination Question Answering (QA), where LLMs are prompted to answer 198 multiple-choice questions from the 4 games for evaluation of three key reasoning abilities: Environment Comprehension, ToM Reasoning, and Joint Planning. Furthermore, to enable LLMs for multi-agent coordination, we introduce a Cognitive Architecture for Coordination (CAC) framework that can easily integrate different LLMs as plug-and-play modules for pure coordination games. Our findings indicate that LLM agents equipped with GPT-4-turbo achieve comparable performance to state-of-the-art reinforcement learning methods in games that require commonsense actions based on the environment. Besides, zero-shot coordination experiments reveal that, unlike RL methods, LLM agents are robust to new unseen partners. However, results on Coordination QA show a large room for improvement in the Theory of Mind reasoning and joint planning abilities of LLMs. The analysis also sheds light on how the ability of LLMs to understand their environment and their partner’s beliefs and intentions plays a part in their ability to plan for coordination. Our code is available at https://github.com/eric-ai-lab/llm_coordination.
1 Introduction
In a wide range of activities, from daily tasks such as cooking to critical operations like rescue efforts, cooperation without mixed intentions is essential. These scenarios are examples of Pure Coordination Games, where all involved parties benefit from choosing strategies that are perfectly aligned, avoiding any conflict of interest. These games require agents to reason about their environment and plan while considering the beliefs and intentions of their partners. Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents. Previous works have explored the use of LLMs for developing collaborative agents, yet the requisite conditions, strengths, and limitations of LLMs in coordination games remain unclear. In this study, we intend to bridge the gap by performing a comprehensive evaluation and analysis of the multi-agent coordination abilities of LLMs.
Therefore, we introduce a new LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1. Agentic Coordination and 2. CoordinationQA. In Agentic Coordination, LLMs are scaffolded with components that allow them to act within actual game environments, providing a holistic evaluation of the competencies of LLMs to act as coordination agents. In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents are required actively cooperate with their partners. The benchmark includes four collaborative games, providing a comprehensive analysis platform.
To enable LLMs for multi-agent coordination, we present a Cognitive Architecture for Coordination (CAC) framework that facilitates LLM interaction with game environments in a plug-and-play approach. CAC translates game elements into textual formats and leverages auxiliary LLMs for improved coordination to enable effective multi-agent collaboration. Our Experiments on the Agentic Coordination task with CAC reveal that Large Language Models are capable of understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games. They exhibit these competencies without any training, fine-tuning, or few-shot examples.
A comparative analysis on the Agentic Coordination task reveals that LLM agents outperform Multi-agent RL solutions in games that require minimum theory-of-mind reasoning and focus on taking commonsense actions based on the environment. However, they struggle in more complex settings where agents need to actively consider their partner’s beliefs and intentions. We also observe that LLM agents are capable of collaborating with new partners, unlike self-play MARL methods (Carroll et al., 2019a; Bard et al., 2020) that fail to adapt to unseen agents.
For a more nuanced analysis of the coordination abilities of LLMs, we create the CoordinationQA Suite. This suite is designed to dissect the capabilities of LLMs in single-turn reasoning within coordination games, focusing on three key areas: Joint Planning, Theory of Mind (ToM), and Environment Comprehension. Joint Planning evaluates LLMs’ decision-making for optimal long-term outcomes, ToM questions probe the understanding of partner agents’ intentions and needs, and Environment Comprehension assesses knowledge of game rules and dynamics. Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types. LLMs are most proficient in Environment Comprehension, indicating they understand game rules and states well. However, they face significant challenges in Theory of Mind Reasoning, with difficulty inferring others’ intentions and needs. This issue worsens in Joint Planning, where most LLMs underperform, some even worse than random choices. These results highlight LLMs’ limited reliability and effectiveness as coordination partners. Additionally, our analysis shows a moderate to strong correlation between Theory of Mind Reasoning, Environment Comprehension, and Joint Planning performance, underscoring their role in effective coordination.
In summary, our contributions are four-fold:
1.
We introduce the LLM-Coordination Benchmark for evaluating and analyzing LLMs in Pure Coordination Games, covering multi-turn Agentic Coordination and single-turn Coordination QA tasks.
2.
We develop the plug-and-play Cognitive Architecture for Coordination Framework to enable LLMs to effectively participate in complex, partially observable coordination games like Hanabi, demonstrating the first zero-shot application for LLM agents in such scenarios.
3.
We perform a holistic evaluation of LLM agents in Self-play and Cross-play settings, offering a detailed comparison with RL baselines and showcasing their potential as Coordination Agents.
4.
We investigate Environment Comprehension and Theory of Mind Reasoning as essential components of LLMs’ overall Joint Planning capabilities, highlighting their critical importance in coordination tasks.
2 LLM-Coordination Benchmark
2.1 Multi-turn Agentic Coordination
In the Multi-turn Agentic Coordination task, LLMs participate in end-to-end pure coordination games as agents, where the best strategy for all participating agents is to cooperate. LLMs under test are plugged into coordination frameworks with components like memory modules and grounding modules to act in complete games. These LLM agents can then be partnered with any policies or agents to complete the games.
Our LLM-Coordination benchmark includes 4 pure coordination games, Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), Collab Capture and Collab Escape.
Hanabi Challenge: In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5). A unique aspect of the game is that the players can only view their partner’s cards, not their own. This requires players to work collaboratively, utilizing reveal tokens to provide hints about the cards in their partner’s hand. These hints can be about either the color or the rank of the cards. For instance, using a single hint token, a player can indicate all cards of a certain rank in their partner’s hand. Hanabi serves as an exemplary Pure Coordination game, necessitating player cooperation to achieve optimal outcomes. Success in Hanabi hinges on the ability to understand partners’ perspectives, navigate decisions based on incomplete information, and engage in implicit communication, making it an excellent testing ground for coordination among agents.
Overcooked-AI: In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups. This environment includes a variety of layouts, each with its own arrangement and quantity of onion dispensers, plate dispensers, cookers, delivery zones, and countertops. To prepare a dish, agents are required to insert three onions into a cooker, initiating a cooking process that lasts 20 time steps. Upon completion, the soup must be plated and delivered to complete the task. Each layout presents unique challenges, emphasizing the need for agents to comprehend their surroundings, locate necessary resources, and synchronize their actions with their teammate for effective collaboration.
Collab Capture: Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms. The rooms are connected by doors, which can be controlled through access buttons that can be found in other rooms. The agents’ task is to capture the adversary in the least amount of time using effective strategies, including cornering the adversary and manipulating the doors to enable their partner or disable the adversary.
Collab Escape: Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms. They need to fix two generators (similar to the game Dead-by-Daylight (Dea, 2016)) located in different rooms to open an exit portal. The adversary tries to catch the agents, and the win condition is any one agent escaping. This game requires strategies like luring the adversary away from the partner, sacrificing for the partner’s safety, and manipulating the movement of the adversary.
2.2 Single-turn Coordination QA
The agentic coordination task paints a holistic picture of the abilities of LLMs as agents. To dive deeper into the specific strengths and weaknesses of LLMs, we develop the CoordinationQA Suite. Inspired by the idea of Unit Testing for evaluating AI agents Knott et al. (2021), we manually sampled edge cases from all 4 pure coordination games mentioned in Section 2.1. All of these edge cases necessitate agents to actively understand their current state, think about their partner’s intentions, and come up with the best plans for coordination. We then create a set of three types of questions for each scenario in our CoordinationQA Suite.
Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment.
Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners.
Joint Planning (JP) questions provide agents with the state/observation and ask them to predict the best next action for effective coordination. This question is essentially the same question that LLMs need to repeatedly solve when they act as agents.
All the questions were manually developed and labeled. We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear optimal solutions. We generated a total of N=66 scenarios (25 from Overcooked, 28 from Hanabi, and 13 from the two Collab Games) and created 3 questions per scenario, resulting in 198 unique questions. The right side of Figure 1 demonstrates the sampling process for the three types of questions with an example from the game Overcooked. The selected scenario shows the Blue agent about to place their third onion in the cooker, and the green agent needs to figure out what to do next.
3 Cognitive Architecture for Coordination
We develop a LLM Agent architecture based on Sumers et al. (2023), which we dub Cognitive Architecture for Coordination (CAC), for multi-agent coordination. Using CAC we can easily plug and play a LLM agent, and pair it with any partner (e.g., another CAC agent, a human player, or other AI agents). The architecture consists of three key elements: Memory, Reasoning, and Grounding.
Memory Module includes (1) Long-Term Memory for storing the Game Description including the game’s rules, conventions, objectives and action space, (2) Working memory which consists a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent. In the example shown in Figure 2 the Long Term Memory includes a description of the game of Hanabi, The Working memory includes observations including the current stack, partner’s hand, beliefs and knowledge of both agents and information regarding the available tokens, and cards, and the Episodic Memory includes the previously selected discards, plays and hints.
Reasoning Module is where the Large Language Model (LLM) is plugged into the framework. It takes the textual description in the working memory as input and generates the next best action based on the context. For coordination games like Hanabi that require a more sophisticated Theory of Mind reasoning, we add an auxiliary Theory of Mind Reasoning LLM whose sole responsibility is to interpret the partner agent’s actions and requirements. This extra reasoning is added to the working memory before passing to the primary LLM. Additionally, we also utilize a Self-verification LLM which verifies the safety of the selected action. In the Hanabi example in Figure 2, the reasoning module interprets the provided clue of ”Revealing Rank 1 Cards” and decides to play one of the pointed cards on the empty stacks. The generated action is passed to the grounding module.
Grounding Module is responsible for interfacing the reasoning and memory modules’ textual decision-making spaces with the actual game mechanics. Its primary task is to take the selected action from the reasoning module and translate it into game-compatible action(s). The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like ”pick up onion from o0.” into sequences of lower-level actions. On the other hand, in games like Hanabi, it just needs to match actions like ”Reveal Bob’s Red Color Cards” to their lower-level representations. The grounding module is also responsible for the secondary task of filtering out infeasible actions based on the context of the game.
4 Experiments
4.1 Agentic Coordination
4.1.1 Setup
We perform two types of experiments in agentic coordination: Self-Play and Cross-Play. In self-play settings, the participating agents are of the same type. In Cross-Play experiments, we pair agents with unseen partners, and they need to adapt their behavior to the actions of these new partners.
Self-play Baselines: For Overcooked we use Proximal Policy Optimization (Schulman et al., 2017) and Population-Based Training (Jaderberg et al., 2017) as baselines for comparison. These baselines were established by Carroll et al. (2019a). The Hanabi challenge has been extensively studied and solved using MARL methods. We use Bayesian Action Decoder (Bard et al., 2020), Simplified Action Decoder (Hu & Foerster, 2021), and Off-Belief Learning (Hu et al., 2021a) as baselines.
Cross-play Baselines: For Overcooked, we use a Behavior Cloning model trained on human data Carroll et al. (2019a) and a Proximal Policy Optimization (PPO) agent trained with the Human Behavior Cloning agent Carroll et al. (2019a) as baselines for comparison. We also use human proxies based on behavior cloning as unseen partners. For Hanabi, we use the Simplified Action Decoder (SAD) as a baseline. We pair our agents with the Off-Belief Learning (Hu et al., 2021a), which was trained to generate grounded policies and adapt to unseen partner agents.
Metrics: We measure the total score achieved by agents in Overcooked, where each delivery provides 20 points to both agents. In the case of Hanabi, the metric is the total number of cards that have been correctly arranged by the players.
4.1.2 Results and Analysis
Overcooked Layouts | |||||
Method | CR | AA | Ring | FC | CC |
PPO (Schulman et al., 2017) | |||||
PBT (Jaderberg et al., 2017) | |||||
CACGPT-4-turbo | |||||
CACGPT-3.5-turbo | |||||
CACMixtral8x7B |
Class | Method | Score |
RL | BAD (Foerster et al., 2019) | |
SAD (Hu & Foerster, 2021) | ||
OBL (Hu et al., 2021a) | ||
CAC | GPT-4-turbo | |
GPT-3.5-turbo | ||
Mixtral-8x7b |
Method | Score |
LLM+Self verif.+ToM | |
LLM+Self Verif. | |
LLM |
LLM Agents outperform or match state-of-the-art RL methods in coordination games that depend more on understanding the environment.
We observed that LLM agents (w. GPT-4-turbo) outperform or match the overall performance of RL methods across all layouts of Overcooked-AI. Table 1 presents the numerical scores attained by different agents when paired with a partner agent of the same type. This implies that LLM agents outdo RL agents that have been trained together through Self-play without any game-specific training or fine-tuning. It is, however, important to note that LLM agents are significantly slower and larger than RL models and are not fit for real-time use yet (latency (seconds) of with Chain-of-thought and without for GPT-4-turbo). Furthermore, other models we tested, GPT-3.5-turbo and Mixtral8x7b, are faster but fall short of the RL baselines. We also see positive results on the CollabCapture and CollabEscape games with CAC agents (w. GPT-4), achieving a 100% success rate. However, other LLMs are unable to crack CollabEscape (see Appendix D).
LLM agents struggle at effective planning when advanced Theory of Mind reasoning is required. In Hanabi Challenge, LLM agents seem to struggle compared to RL methods (see Table 2). Among all LLMs, GPT-4-turbo performs reasonably well, while other LLMs can barely complete the Hanabi games. We attribute this failure to two factors. First, there is little room for errors in Hanabi. Any misplay or mis-clue leads to the loss of a life token. Second, Hanabi requires much more complex Theory of Mind Reasoning compared to the Overcoked-AI environment. Each action requires agents to actively consider their partner’s beliefs, intentions, and how they would react to implicit communication.
In contrast, Overcooked is fully observable, and its action space consists of actions like pick up an onion from onion_dispenser_0 and place onion in cooker_0. Under most scenarios and layouts, LLMs only need to consider the next best steps based on the state of the environment. For example, We conduct an ablation study of removing partner inventory and location, which reveals minimum impact on overall performance (1 less delivery in Cramped Room and Asymmetric Advantage layouts each in 100 timesteps), showing that the primary challenge for LLMs in games like Overcooked is the Environment Comprehension ability.
LLM agents benefit from auxiliary reasoning modules in imperfection information games with low room for errors. Without the support of auxiliary modules, LLM agents seem to bomb (lose all three lives) in every game (see Table 3). To rectify this, our CAC framework incorporates auxiliary LLM-powered modules, including an Explicit Theory of Mind Reasoning (ToM) module and an Answer Verification (AV) module. The answer verification module is a game-changer in its ability to stop LLM hallucinations about environmental facts from causing fatal mistakes, thus reducing the chance of mis-plays. The ToM reasoning LLM delegates the responsibility of interpreting partner clues and understanding partner needs to different LLMs, allowing the primary LLM to focus on synthesizing the available information to plan the next action.
Overcooked Layouts | |||||
Method | CR | AA | Ring | FC | CC |
BC (Carroll et al., 2019a) | | | | | | | | | | |
PPOBC (Schulman et al., 2017) | | | | | | | | | | |
CACGPT-4-turbo111For CAC, we run a single trial from either position due to cost constraints. In Table 1 we have observed that the performance of CAC agents does not vary by more than one delivery. | | | | | | | | | | |
Method | Self-Play | Cross-Play w/ OBL-1 | Cross-Play w/ OBL-4 |
SAD (Hu & Foerster, 2021) | |||
CACGPT-4-turbo |
LLM Agents are robust to unseen partners. We use Overcooked-AI and the Hanabi challenge as testbeds to evaluate the performance of LLM agents when paired with unseen agents. This task is popularly known as Zero Shot Coordination. For experiments in Overcooked-AI, we pair our LLM agents as well as baselines with proxy-human agents. These proxy human agents are behavior cloning agents trained using human data by Carroll et al. (2019b). As shown in Table 4, we discover that LLM agents outperform both Behavior Cloning as well as PPO agents trained with human data.
For experiments in Hanabi, we pair our agents with Off-Belief Learning (OBL) agents (Hu et al., 2021a). OBL is a MARL strategy that generates grounded clues and actions and is the state-of-the-art method for both self-play and cross-play in Hanabi. OBL agents provide observation-grounded clues and collaborate well with humans. Therefore, we use them as unseen partners in our experiments. Table 5 shows that CAC agents score an average of 15.00 points with the OBL-1 agent compared to their self-play scores of 13.66. This indicates no degradation in coordination abilities with a new partner. The baseline RL method, Simplified Action Decoder (SAD) Hu & Foerster (2021), fails critically when paired with unseen OBL agents, even though it excels at self-play (22.00 points) due to self-play training.
MARL agents trained with self-play struggle when paired with unseen partners in common payoff tasks, because they converge to arbitrary policies that only the two partners involved in the self-play training understand (Carroll et al., 2019a; Bard et al., 2020). Since LLM agents haven’t been explicitly trained to play these games, they base their outputs on the provided textual observation and commonsense knowledge learned from pre-training, and thus are much more robust to unseen partners.
4.2 Coordination QA
Setup.
We assess the performance of 6 Families of Large Language Models (LLMs) Jiang et al. (2023; 2024); Touvron et al. (2023); Chiang et al. (2023); OpenAI (2023) across three dimensions: Environment Comprehension (EC), Theory of Mind Reasoning (ToM), and Joint Planning (JP). For each category, LLMs respond to multiple-choice questions (MCQs), with their responses evaluated against ground-truth answers through fuzzy string matching. To account for the variability in LLM responses, we conduct three trials per model. We also report a Random baseline.
Comparative Results of LLMs in Environment Comprehension, ToM Reasoning, vs. Joint Planning. In Figure 3, we see that most LLMs achieve their best results on the Environment Comprehension question. The best performing LLM GPT-4-turbo gets more than 80% Environment Comprehension Questions correct. The overall performance across LLMs drops on the more challenging Theory of Mind reasoning questions, but GPT-4-turbo is still competent, reaching a 54% accuracy. The overall accuracy of LLMs on Joint Planning questions is still significantly weak, with even the best LLM scoring less than 40%, indicating a large room for improvement in LLMs’ ability to perform coordination reasoning. Another cause for concern is that open-source LLMs perform abysmally at Joint Planning, with some models performing worse than a random baseline.
Variables | ||
ToM | 0.813 | 0.389 |
EC | 0.895 | 0.506 |
Impact of Environment Comprehension and ToM Reasoning abilities on Joint Planning. Having defined Joint Planning as the capacity of an agent to select the appropriate subsequent action based on available information, we argue that proficiency in Environment Comprehension and Theory of Mind Reasoning is crucial for adept Joint Planning, and LLMs that do well at these two will do well at JP. Correlation analysis of ToM and EC with JP across the data from Figure 6 reveals that ToM has a moderate positive correlation to JP, whereas EC shows a strong positive correlation.
5 Related Work
Multi-agent Coordination
In Game Theory, Pure Coordination games are situations where the payoff is commonly shared between all agents. In such situations, cooperating is the best strategy. Various benchmarks and games have been used to evaluate Multi-Agent Coordination abilities over the years including Multiparticle Environment Lowe et al. (2017), Overcooked-AI Carroll et al. (2019a), and the Hanabi Challenge Bard et al. (2020). The foundational work by Carroll et al. (2019a) emphasized the significance of incorporating human data for effective human-ai collaboration. Subsequent research on the Overcooked-AI challenge has pivoted towards enabling self-play-trained agents to coordinate seamlessly with humans within this environment. These studies employ various techniques, including self-play with past agent checkpoints (Strouse et al., 2021), centralized population entropy objectives (Zhao et al., 2023), open-ended objectives using graph theory (Li et al., 2023a), policy ensembles with context-aware mechanisms (Lou et al., 2023), and the incorporation of human biases as linear hidden rewards (Yu et al., 2023), to enhance the training and diversity of AI agents in different scenarios. On the Hanabi Challenge much effort has been made to learn grounded policies Hu et al. (2021b; a) over arbitrary conventions. Embodied environments usually set up in household environments have also been recently used to study multi-agent coordination (Puig et al., 2021; Jain et al., 2020; 2019; Gan et al., 2021). The Overwhelming majority of approaches to coordination problems have focused on utilizing and enhancing Reinforcement Learning methods to solve the problems of multi-agent coordination. In this work, we argue that Large Language Models are an alternative approach to these coordination problems as they show emergent reasoning abilities, demonstrate theory-of-mind-like abilities, and do not converge to policies that cause arbitrary joint interactions.
Planning and Reasoning with Large Language Models
Large Language Models (LLMs) have demonstrated remarkable capabilities of reasoning in natural language (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art performance across a spectrum of verbal reasoning tasks. It was then shown that LLMs could be augmented with components like memory, tools, perception, and grounding to create agents that could interact with an external environment (the web, simulators, games, etc.) These LLM agents have shown to be capable of solving long-horizon tasks, playing complex games (Wu et al., 2023; Wang et al., 2023) and interacting with simulated embodied environments (Liang et al., 2022; Song et al., 2022). Zhang et al. (2023b) developed a modular agent framework that was capable of cooperating with partner agents in embodied spatial rearrangement problems, demonstrating increased efficiency through coordination. Zhang et al. (2023a) develop a specialized architecture that enables LLMs to play in the Overcooked-AI environment. Li et al. (2023b) evaluate and show emergent collaborative abilities of LLMs in gamified simulations. In contrast to existing works, our work focuses on evaluating language agents in established pure coordination games where coordination is not an optional efficiency enhancer but rather a necessity.
6 Conclusion
In this study, we evaluated and analyzed the current large language models in terms of their ability to reason and act in pure coordination games. We introduced the LLM-Coordination benchmark with its two tasks: 1. Agentic Coordination and 2. CoordinationQA. These settings allowed us to conduct holistic comparative studies of LLMs as agents as well as dive deeper into the fine-grained aspects of LLMs as coordination reasoners. We juxtaposed LLM agents with existing Multi-agent Reinforcement Learning agents, discussing the conditions in which LLMs thrive and fail. Finally, we discussed the Theory of Mind Reasoning and Environment Comprehension as prerequisites for coordination and evaluated existing LLMs on these two components.
Appendix A Overcooked Implementation Details
A.1 Game and Layout Description
We use a general game description that explains the rules and objectives of overcooked. Since each layout has a different number of locations, like onion dispensers and cookers, we include a succinct description of each environment , which includes how many instances of particular facilities there are. For environments that include partitions, we mention which partition each of the agents is situated in and what facilities that agents can access. In addition, we also mentioned the shape of the environment.
A.2 State Description
The State is represented in natural language in the working memory, which can be processed by a Large Language Model (LLM). The state includes variables that fully represent the necessary details of the layout as well as the players. The information provided in is equivalent to what would be accessible to a Reinforcement Learning (RL) agent in the form of state representations. The following information is included in :
Objects Held by Each Player
The state description begins by detailing the inventories and of Alice and Bob, respectively. Each inventory (where ) can contain one of the following items: {”onion”, ”plate”, ”cooked soup”}. This inventory information is translated into natural language and incorporated into in the format: “I am holding . Bob is holding .” Such information is vital for inferring the likely subsequent actions of the partner agent.
Location of the Agent Controlled by LLM:
Given the limitations of Large Language Models (LLMs) in interpreting grid-based spatial information, we opt to provide processed location data to the LLM. For each agent (where ), and for each location of interest denoted as loc, we calculate the distance as the number of steps required to reach loc from using the shortest available path. The state description then includes this processed location information in the format: “loc is units away.” Here, loc can represent various points of interest such as onion dispensers, plate dispensers, cookers, delivery areas, kitchen counters, or shared counters. If a location is either inaccessible or blocked by another agent, this is explicitly stated in . For example, if a location is blocked by Bob, it would be stated as “loc is blocked by Bob.” To distinguish between the location information relevant to each agent, prefixes the respective sections with “Your location information:” for the agent controlled by the LLM and “Bob’s location information:” for the partner agent.
Cooker Information
The state description also incorporates information about the cooker, which is central to the gameplay strategy. Specifically, for each cooker , includes the number of onions currently in the pot. Additionally, provides the operational state of the cooker, denoted as , which can be either ”Off” or ”On”. Lastly, the current condition of the soup in the cooker is represented by , which can take one of the following values: ”Cooking”, ”Cooked”, or ”Not Started”. Thus, the information for cooker is formatted as: “ has onions. is . Soup in is .”
Kitchen Counter Information
The state description includes information about kitchen counters, which are primarily used for temporary object storage. Specifically, identifies the closest empty kitchen counter and the set of all counters currently holding an object.
Shared Counter Information
Shared counters serve as specialized kitchen counters for object transfer between agents. For each shared counter , includes the status for , as “ is empty” or “ contains onion,” to offer a complete environmental overview. Unlike kitchen counters, where only the closest empty counter is mentioned, all empty shared counters are mentioned.
Appendix B Hanabi Implementation Details
B.1 Game Description
We structure the game description of Hanabi into the overall objective, the rules of the game, and list of conventions based on the H-group conventions han (2024). We do not use advanced conventions like Chop Cards but stick to basic conventions about the card layout, play clues, and save clues.
B.2 State Description
The state description includes the current Stack , the player’s knowledge of their cards (updated based on clues), the partner agent’s cards , the partner agent’s knowledge of their cards (updated based on previous clues), each card in the discard pile , the remaining Life Tokens , and reveal tokens and the remaining Deck Size . We also precalculate the next card that goes on each stack since LLMs frequently fail to count which card should go next on each stack.
Appendix C Examples of prompts of LLMs used in CAC framework
C.1 Action Generator
The action generator generates a brief explanation and selected the next action to be played.
C.2 Theory of Mind Reasoning LLM
The Theory of Mind Reasoning LLM interprets partner actions and clarifies partner requirements.
C.3 Self Verification LLM
The prompt for the Self Verification LLM is provided as a system prompt. We observed that LLMs were more likely to act as strict verifiers when the prompt is provided as a system prompt.
Appendix D Results of Different LLMs on CollabCapture and CollabEscape
Table 7 summarizes the performance of the three LLMs GPT-4-turbo, GPT-3.5-turbo and Mixtral-8x7b on CollabGames - Collab Capture and Collab Escape. GPT-4-turbo achieves a 100% success rate in both games. In general, CollabEscape was a more difficult game to solve than CollabGames since agents needed to perform sacrificial moves and lure the killer away, requiring more advanced Theory of Mind inference.
Collab Capture | Collab Escape | |||
CAC model | Capture Rate | Turns to Capture | Escape Rate | Turns to Escape |
GPT-4-turbo | 1.00 | 3.99 | 1.00 | 24.67 |
GPT-3.5-turbo | 0.50 | 8.49 | 0.00 | N.A. |
Mixtral-8x7b | 0.75 | 3.88 | 0.00 | N.A. |
Greedy Baseline | 0.50 | 6.00 | 0.00 | N.A. |
Appendix E Generating Questions for CoordinationQA
E.1 Environment Comprehension Questions
The Environment Comprehension (EC) questions are indirect formulations regarding spatial aspects of the layout. In order for an agent to correctly answer an EC question, they must have an understanding of the dynamic details of the current state, the rules of the game, and exhibit spatial awareness. As such, when creating the EC questions, we carefully comb through a given scenario in search of salient points to probe an agent’s understanding of the given environment. Some examples include:
E.2 Theory of Mind Reasoning Questions
There are two primary question types in Hanabi for ToM Reasoning questions. In the first type, we ask the LLM about what information the partner agent needs, while in the second type, we ask it to make inferences about the partner agent’s last action. For all games apart from Hanabi, the ToM questions ask the models to predict the next intended action of the partner agent
E.2.1 Hanabi Question Type-1
E.2.2 Hanabi Question Type-2
E.2.3 Other Games
E.3 Joint Planning Questions
Joint planning questions are effectively the same questions that the LLM solves when they are part of an agentic framework. For each scenario, we ask the LLM to answer the question: ”What is the best next action?”.